OpenStudy (anonymous):

identify the POWER in the radical below. ^7square{3^4}

OpenStudy (across):

I am sorry, but I cannot understand your notation. Could you re-write it? :/

OpenStudy (anonymous):

yes sorry :/

OpenStudy (anonymous):



OpenStudy (anonymous):

The square root sign power is always understood to be two when there is no power actually stated.\[7\sqrt[2]3^{4}\] So you don't normally ever see it written.

OpenStudy (across):

It is important to realize that \(\sqrt{x}=(x)^{1/2}\). Therefore, \(7\sqrt{3^4}=7(3^{4/2})=7(3^2)\).